Circular No. 7427

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1118+480 C. A. Haswell, Open University; D. Skillman, Laurel, MD; J. Patter-

son, Columbia University; R. I. Hynes, Southampton University; and W. Cui, Massachusetts Institute of Technology, on behalf of a larger collaboration, report: “Hubble Space Telescope observed XTE J1118+480 for a total integration time of 2.54 hr (range 115–170 nm) on Apr. 8. The power density spectrum (PDS) is roughly flat below ∼ 0.02 Hz and breaks into a power law (roughly 1/f) above. Such a PDS shape is characteristic of blackhole candidates or ‘atoll-type’ low-mass x-ray binaries in the low state. In addition, the ultraviolet PDS reveals a broad quasiperiodic oscillation (QPO) at 0.08 Hz with FWHM 0.02 Hz. Unfiltered CCD observations (0.68-m f/2.9 telescope) during Apr. 30–May 3 with 0.32-Hz sampling show a QPO at 0.101 Hz (FWHM 0.008 Hz). Both the low-state PDS shape and the QPO are confirmed by preliminary results from RXTE, which indicate that the source has remained in such a low state throughout the outburst. Comparison of simultaneous HST near-ultraviolet (160–235 nm) and RXTE lightcurves reveals a correlation on timescales of seconds. A preliminary analysis indicates that ultraviolet variability lags behind x-rays by 1–2 s. This is likely due to light-echoes caused by reprocessing of the x-ray flares within the binary system; the delay is consistent with expected light-travel-time delays within a 4-hr x-ray binary (IAUC 7397). Further high-time-resolution observations at all wavelengths are urged to investigate light-

echoes further (see http://phys-ftp.open.ac.uk/pub/1118/satobs).”

K. Yamaoka, Y. Ueda, and T. Dotani, Institute of Space and Astronautical Science; and P. Durouchoux and J. Rodriguez, Service d’Astrophysique, Saclay, report observations of XTE J1118+480 with the ASCA satellite during from May 11.527–12.031 UT: “The x-ray intensity showed rapid variability over an averaged flux level of 8 × 10^{−10} erg cm^{−2} s^{−1} (2–10 keV). A QPO feature was found at 0.115 ± 0.002 Hz with an amplitude of ∼ 6 percent rms in the averaged power spectrum; this QPO was confirmed by simultaneous RXTE data, which covered two epochs during May 11.729–11.748 and 11.781–11.931. The QPO frequency has shifted from the value reported before May 4 (0.085 ± 0.002 Hz) by Revnivtsev et al.

(http://xxx.lanl.gov/abs/astro-ph/?0005212). The ASCA GIS spec-

trum in the range 0.7–10 keV is ∼ represented by a power law, over which a slight soft excess is seen below 2 keV. Applying a disk blackbody and power-law model yields an acceptable fit with a temperature of the innermost disk of 0.2 ± 0.1 keV, and a power-law photon index of 1.76 ± 0.02, with an upper limit to an absorption column density of 10^{21} cm^{−2}.”

2000 May 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green